Solve the following recurrence using the master method:

1))2, with T(0) = 2 T(n) (T(n
Answer #1
T(n) = (T(n-1))2 and T(0) = 2
This is not in the form of T (n) aT( ) + f(n) so, we can not masters theorem here
using substitution method:
-----------------------------
T(n)
= (T(n-1))2
= (T(n-2))4
= (T(n-3))8
= (T(n-3))2^3
...
= (T(n-n))2^n
= (T(0))2^n
= 22^n
Answer: T(n) = O(22^n)